A circuit with a power supply of 12V, a capacitor, and a resistor is assembled. The capacitance of the capacitor is 4.7µF. The time it takes for the potential difference across the resistor and the capacitor to become equal is 43 seconds, and the RC time constant is 60s.

Calculate the value of the resistance


1
Expert's answer
2021-09-03T18:45:07-0400

Since "RC = 60s" and "C = 4.7 \mu F = 4.7\times 10^{-6}F", the resistance can be found as follows:


"R = \dfrac{60s}{C} = \dfrac{60s}{4.7\times 10^{-6}F} \approx 13\times 10^6\Omega = 13\space M\Omega"

Answer. "13\space M\Omega".